OpenStudy (anonymous):

Find the equation of the line that passes through (2, -7) and is parallel to the line that passes through (-4, 4) and (-6, 6)

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

Need some help?

OpenStudy (anonymous):

Yes, very lol

OpenStudy (anonymous):

Alright so what do you have so far?

OpenStudy (anonymous):

Just the problem. I'm usually pretty good at math, but I don't get how to do this problem at all.

OpenStudy (anonymous):

It's quite alright.. So we need to find the equation for a line parallel to the one that passes through the two points but also has to pass through a specific point correct? So do you know what type of equation we need? I mean slope intercept, standard, or point slope?

OpenStudy (anonymous):

It just says equation for the problem, and parallel. Which one do you think would work best for that?

OpenStudy (anonymous):

well i would say point slope but if not we could easily change the form when we are done... sooo... shall we begin? haha

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

okay so to start we need to find the equation for the line that passes through, (-4,4) and (-6,6).. do you know how to do that?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

alriight so to do that the easiest would be to use point slope form.. do you know the formula for point slope form?

OpenStudy (anonymous):

no lol, im sorry

OpenStudy (anonymous):

its quite alright :) so the formula is: \[y _{2}-y _{1}=m(x _{2}-x _{1})\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yes, so if we label our points with the x and y intercepts, we can plug them into the formula... so we get: (x1,y1) ( -4, 4 ) and... (x2,y2) (-6, 6 )

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

try plugging them into the equation.. let me know what you get

OpenStudy (anonymous):

6 - 4 = m(-6 - -4)

OpenStudy (anonymous):

great job :D but remember the two negatives turn to a positive sooo... 6-4=m(-6+4) so now we find the slope, or m..

OpenStudy (anonymous):

0 = m

OpenStudy (anonymous):

or 12 = m

OpenStudy (anonymous):

\[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1}}\]

OpenStudy (anonymous):

m = 6 - 4 / -6 - 4

OpenStudy (anonymous):

close remember the negative turns positive on the bottm so -6+4

OpenStudy (anonymous):

ok m = 6 - 4 / -6+ 4

OpenStudy (anonymous):

yes, now simplify :)

OpenStudy (anonymous):

m = -1/ -1 m = 1

OpenStudy (anonymous):

Ooooo... so close :) nice try.. so we have 6-4=2 and -6+4=-2 so 2/-2 or -1

OpenStudy (anonymous):

follow me?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so plug that in for m, the equation for the line would be?

OpenStudy (anonymous):

6 - 4 = -1(-6 - -4)

OpenStudy (anonymous):

-6+4 remember? haha some reason I hear myself like I'm coming off mean or irritating haha

OpenStudy (anonymous):

It's ok haha, so 6 - 4 = -1(-6 + 4)

OpenStudy (anonymous):

yes so to find the equation of the line that is parallel to this one, we simply take the slope and plug the point into a fresh equation and plug in the point

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait is this a multiple choice question? just wondering

OpenStudy (anonymous):

no

OpenStudy (anonymous):

alright... so we get y+7=-1(x-2)

OpenStudy (anonymous):

and we're DONE!!! :DD yay!!!

OpenStudy (anonymous):

so y+7=-1(x-2) is the answer?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

no problem thanks for baring with me haha